one third of a number is 5 less than half of the same number what is that number show all steps i need to get my diploma
Answer by algebrapro18(206)

one third of a number is 5 less than half of the same number

writing this as an algebraic expression you get

1/3x=5-1/2x

to solve this equation you need to get a common denominator. Doing so you get

2x=30-3x

now combine like terms

5x=30

finally divide by 5

x=6

checking

1/3(6)=5-1/2(6)
2=5-3
2=2